OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x2 + 9x + 20 2, 10 −4, −5 −2, −10 4, 5

OpenStudy (cwrw238):

first factor the function into the form (x + a)(x + b) = 0 now find the numbers a and b a * b must = 20 and a + b must = 9

OpenStudy (cwrw238):

lets guess 2 numbers which multiply together to give 20 4 and 5 right? and what do they add up to ?